# Kolmogorov-Arnold Energy Models: Fast, Interpretable Generative Modeling

### v9

@@ Abstract
Generative models typically rely on either simple latent priors (e.g., Variational Autoencoders, VAEs), which are efficient but limited, or highly expressive iterative samplers (e.g., Diffusion and Energy-based Models), which are costly and opaque. We introduce the Kolmogorov-Arnold Energy Model (KAEM) to bridge this trade-off and provide a new avenue for latent-space interpretability. Based on a novel interpretation of the Kolmogorov-Arnold Representation Theorem, which provides new structural biases for generative modeling, KAEM imposes a univariate latent structure that enables fast and exact inference via the inverse transform method. With a low-dimensional latent space and appropriate inductive biases, we show that importance sampling becomes a viable, unbiased, and highly efficient posterior inference method. For settings where importance sampling fails, we propose a population-based strategy that decomposes the posterior into a sequence of annealed distributions to improve mixing during sampling, a common pitfall in Energy-based Models. We evaluate KAEM on SVHN and CelebA, demonstrating potential for promising sample quality when correct inductive biases are chosen, while improving interpretability and inference speed against VAEs.
@@ 4.1 Setup
We subsequently evaluate the mixture EBM prior against SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015), sampled via component-wise ITS and paired with CNN-decoder architectures, which are slightly smaller variants of those used by Pang et al. (2020). We additionally compare KAEM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the thermodynamic criterion. These comparisons assess the practicality of using mixture densities for scaling KAEM and using annealing to improve training for high-dimensional RGB images. Hyperparameters are provided in Table 5. Architectural details for KAEM and the baseline VAE are provided under Appendix F.2.
@@ 4.2 Results
A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator. Full architectural details are provided under Appendix 3.1. Posterior sampling was conducted with 40 steps of ULA, and a standard Gaussian reference prior was used, $\pi_{0}=\mathcal{N}(0,1)$. KAEM was trained for 200 epochs with a batch size of $N_{x}=100$ on 20,000 training examples. All remaining hyperparameters are in Tab. 5. Annealing was conducted with $N_{t}=5$ power posteriors and the $p$ scheduling outlined in Sec. D.1.
@@
FID KID
@@
FID KID
@@
MLE 251.64 0.176 0.949 0.2687 0.00037 0.418
Thermo 165.62 0.152 0.962 0.1815 0.00021 0.019
VAE 113.33 0.113 0.974 0.1152 0.00013 0.591
Figure 4: Sampling (inference) times of KAEM and VAE across latent dimensionality, $Q$. KAEM uses inverse transform sampling (ITS) from the EBM prior, while VAE decodes from $\bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Error bars correspond to standard deviation. We believe that the VAE is slower because it includes a dense projection layer before the convolutional decoder, while KAEM’s generator operates directly on the reshaped latent representation (see F.2).
We note that sample quality is sensitive to the prior architecture. In Appendix H, we report a smaller-scale experiment using RBF bases instead of wavelets, which resulted in higher-quality samples after just 16,000 parameter updates. This architecture is more promising against VAEs
@@
Our small mixture prior model was able to generate low-quality samples on both SVHN and CelebA, demonstrating the potential of component-wise ITS for exact and highly efficient prior sampling, as an alternative to the iterative and biased approach of Pang et al. (2020). Full computational benchmarks for the architectures and hyperparameters for SVHN ($32\times 32$) are provided in Appendix A, demonstrating that KAEM samples faster than the baseline VAE, (despite requiring more memory). We believe this is due to the slight architectural differences noted in Fig. 4.
Our proposed thermodynamic strategy led to worse image metrics on SVHN, (and both datasets in Appendix H), but enabled KAEM to escape posterior collapse to a single mode when trained on CelebA. As shown in Appendix G.2, training with ULA and MLE produced degenerate CelebA samples, whereas population-based methods successfully mitigated this failure mode.
Subjectively, images generated under thermodynamic training appeared more well-defined (i.e., lower variance), but were less diverse than under MLE. Future work could investigate prior parameterized by functions other than wavelets, as well as the regimes in which thermodynamic training is preferable to MLE. While SVHN performed better under the MLE objective, CelebA appeared to benefit subjectively from thermodynamic training, producing more human-perceptible features in both experiments, wavelets and RBFs. For instance, a better tuned annealing schedule may sharpen SVHN features similar to CelebA in Fig.18.
@@
Ultimately, we cannot justify the thermodynamic strategy due to its substantial computational overhead. With greater VRAM availability, it would be possible to parallelize a larger number of temperatures than the $N_{t}=5$ used in this study, which could potentially improve sample quality, as suggested by the results of Calderhead and Girolami (2009).
Alternatively, different posterior sampling approaches, such as other annealing schedules, more ULA steps, or the use of autoMALA (Biron-Lattes et al., 2024) with bias correction and temperature-adaptive step-size tuning, may further improve performance. However, exploring these options might not be worthwhile until the hardware described by Zettascale Computing Corporation (2026) becomes available, given the added algorithmic complexity, hyperparameter sensitivity, and extended training times involved.
Another direction for posterior inference could involve amortized approaches similar to Kingma and Welling (2022), particularly since the VAE consistently outperformed KAEM. The information bottleneck criterion has previously been derived for latent EBMs by Kong et al. (2021).
@@ 5 Conclusions
KAEM provides a principled framework for structured and interpretable generative modeling. Our results demonstrate that importance sampling enables efficient training on simple or low-dimensional datasets, while the Kolmogorov-Arnold Representation Theorem offers a principled basis for architectural design, allowing KAEM’s structure to be determined entirely by its latent space dimensionality. The inverse transform method provides clear advantages over Langevin Monte Carlo for sampling energy-based priors, offering both speed and exactness. While the benefits of annealing and population-based posterior sampling remain inconclusive in our experiments, these directions warrant further investigation as hardware capabilities continue to advance, particularly with developments such as those introduced by Zettascale Computing Corporation (2026).
@@ 5.1 Future work
Moreover, while ITS enables fast and scalable ismaller nference for mixture priors, KAEM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of Langevin dynamics. GPUs are poorly suited to Langevin methods because of irregular memory access patterns and bandwidth pressure from differentiating with respect to inputs, which requires storing intermediate outputs across generator layers.
@@
The smaller experiment in Appendix H demonstrated that the choice of prior architecture and its encoded inductive biases significantly affect sample quality—the alternative basis achieved superior image metrics after only 16,000 parameter updates. Future work could further investigate this phenomenon across different KAN bases and datasets, as well as explore how domain knowledge can guide the selection of univariate functions.
@@ F.1 Hyperparameters tables
grid_update_ratio $\alpha_{\text{grid}}$ N/A
Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1.5,1.5][-1.5,1.5]
@@
generator_noise $\sigma_{\epsilon}$ 0.25
generator_variance $\sigma_{\epsilon}$ 0.25
@@
update_prior_grid (-) false
update_llhood_grid (-) false
num_grid_updating_samples (-) 100
grid_update_frequency (-) 100
grid_update_decay (-) 0.999
@@
Number of epochs $N_{\text{epochs}}$ 200
eps $\epsilon$ 0.001
@@ F.2.1 KAEM Architectures
Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
@@ 4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
Config: kernel=4,4,4,4 strides=1/2/2/2, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
@@
Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
1 $\times4$ convT($512$), LeakyReLU $4\times 4\times 512$ 1
2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT(64), LeakyReLU $32\times 32\times 64$ 2
@@ 5 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
Config: kernel=4,4,4,4,4 strides=1/2/2/2/2, paddings=0/1/1/1/1, activation=LeakyReLU, batchnorm=False
@@ F.2.2 Baseline VAE architectures
1 $\times4$ conv($64$), LeakyReLU $16\times 16\times 64$ 2
2 $\times4$ conv($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ conv($256$), LeakyReLU $4\times 4\times 256$ 2
4 Flatten + FC $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
Config: kernels=4,$4,$ strides=2/2/2, paddings=1/1/1, latent_dim=81, batchnorm=False
@@ 4 Flatten + FC $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
1 FC + Reshape $4\times 4\times 256$
2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
@@ 4 $\times4$ convT(33), Sigmoid $32\times 32\times 3$ 2
Config: kernels=4,$4,$ strides=2/2/2, paddings=1/1/1, latent_dim=81, batchnorm=False
@@
1 $\times4$ conv($64$), LeakyReLU $32\times 32\times 64$ 2
2 $\times4$ conv($128$), LeakyReLU $16\times 16\times 128$ 2
3 $\times4$ conv($256$), LeakyReLU $8\times 8\times 256$ 2
4 $\times4$ conv($512$), LeakyReLU $4\times 4\times 512$ 2
5 Flatten + FC $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
Config: kernels=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, latent_dim=81, batchnorm=False
@@ 5 Flatten + FC $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
1 FC + Reshape $4\times 4\times 512$
2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT($64$), LeakyReLU $32\times 32\times 64$ 2
@@ 5 $\times4$ convT(33), Sigmoid $64\times 64\times 3$ 2
Config: kernels=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, latent_dim=81, batchnorm=False
@@ G.2 SVHN & CelebA
Appendix H Smaller RBF Experiment
To investigate how hyperparameter choice affects sample quality, we conducted a smaller-scale supplementary experiment using RBF bases. Most hyperparameters were kept consistent with Tab. 5, except for using RBF bases with ReLU base activation and 20 grid points, alongside a smaller additive noise variance of $\sigma_{\text{noise}}=0.1$ and a larger posterior sampling step size, $\eta=0.01$. Training was conducted for 40 epochs with 20,000 training examples and a batch size of $N=50$ (i.e., only 16,000 parameter updates).
The grid updating procedure proposed by Liu et al. (2024) was applied periodically, starting at once every 100 parameter updates with a frequency decay of 0.999, using 100 samples to update the prior’s domain based on regions explored by ULA.
Effectively unbiased metrics were collected using regression on sample sizes 1000, 1200, 1400, 1600, 1800, 2000, 2200, 2400, 2600, 2800, 3000 (i.e., significantly smaller than those used in the main experiment). Regression statistics, (standard errors and $R^{2}$), were not collected for this supplementary experiment.
Model $\overline{\text{FID}}_{\infty}$ $\overline{\text{KID}}_{\infty}$
MLE 78.41 0.0627
Thermo 165.79 0.1608
Table 13: SVHN ($32\times 32$), RBF experiment.
Model $\overline{\text{FID}}_{\infty}$ $\overline{\text{KID}}_{\infty}$
MLE 130.05 0.1427
Thermo 150.25 0.1637
Table 14: CelebA ($64\times 64$), RBF experiment.
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 17: Generated SVHN ($32\times 32$) samples using RBF bases with mixture prior architecture
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 18: Generated CelebA ($64\times 64$) samples using RBF bases with mixture prior architecture

### v10

@@ Abstract
Generative models typically rely on either simple latent priors (e.g., Variational Autoencoders, VAEs), which are efficient but limited, or highly expressive iterative samplers (e.g., Diffusion and Energy-based Models), which are costly and opaque. We introduce the Kolmogorov-Arnold Energy Model (KAEM) to bridge this trade-off and provide a new avenue for latent-space interpretability. Based on a novel interpretation of the Kolmogorov-Arnold Representation Theorem, which provides new structural biases for generative modeling, KAEM imposes a univariate latent structure that enables fast and exact inference via the inverse transform method. With a low-dimensional latent space and appropriate inductive biases, we show that importance sampling becomes a viable, unbiased, and highly efficient posterior inference method. For settings where importance sampling fails, we propose a population-based strategy that decomposes the posterior into a sequence of annealed distributions to improve mixing during sampling, a common pitfall in Energy-based Models. We present initial comparisons of KAEM against VAEs for the SVHN and CelebA datasets, demonstrating its potential for competitive sample quality, inference speed, and interpretability.
@@ 5.1 Setup
We subsequently evaluate the mixture EBM prior against SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015), sampled via component-wise ITS and paired with CNN-decoder architectures, which were smaller and shallower variants of those used by Pang et al. (2020), except with an additional projection layer to ensure exact matches against the baseline VAE decoders they were compared against. Depths and sizes were typically limited by the VAEs, since the additonal encoder network required substantial GPU VRAM.
We also compare KAEM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the thermodynamic criterion. These comparisons assess the practicality of using mixture densities for scaling KAEM and using annealing to improve training for high-dimensional RGB images. Hyperparameters are provided in Table 5. Architectural details for KAEM and the baseline VAE are provided under Appendix F.2.
@@ 5.2 Results
A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator. Full architectural details are provided under Appendix 4.1. Posterior sampling was conducted with 40 steps of ULA, and a standard Gaussian reference prior was used, $\pi_{0}=\mathcal{N}(0,1)$. KAEM was trained for 100 epochs with a batch size of $N_{x}=50$ on 20,000 training examples. All remaining hyperparameters are in Tab. 5. Annealing was conducted with $N_{t}=5$ power posteriors and the $p$ scheduling outlined in Sec. D.1.
@@
$\overline{\textbf{FID}}_{\infty}$ $\overline{\textbf{KID}}_{\infty}$
@@
$\overline{\textbf{FID}}_{\infty}$ $\overline{\textbf{KID}}_{\infty}$
@@
MLE 161.39 0.195 0.907 0.1782 0.00029 0.478
Thermo 174.32 0.096 0.976 0.2053 0.00020 0.299
VAE 131.65 0.143 0.954 0.1403 0.00022 0.281
Figure 4: Sampling (inference) times of KAEM and VAE across latent size, $Q$. KAEM uses inverse transform sampling (ITS) from the EBM prior, while VAE decodes from $\bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Error bars correspond to standard deviation. Given that the decoder/generator architectures were exactly matched, we believe the VAE was faster since it simply decodes from Gaussian noise. KAEM uses ITS from the prior with 200 quadrature nodes. This discrepancy can be reduced without much change in sample quality by reducing the number of quadrature points used, or by optimizing the ITS computational kernel. At its maximum ($Q=101$) KAEM took about $1.7\times$ or 7 milliseconds longer to sample.
@@
Our small mixture prior model was able to generate recognizable samples on both SVHN and CelebA, demonstrating the potential of component-wise ITS for exact and efficient prior sampling as an alternative to the iterative and biased approach of Pang et al. (2020). Full benchmarks are provided in Appendix A. KAEM with ULA is slower to train than VAEs, though importance sampling achieves comparable training times. Sampling is also slower than VAEs, which we attribute to ITS. Nonetheless, KAEM remains faster than iterative EBM samplers and diffusion models.
On both SVHN and CelebA, VAEs achieved the best $\overline{\text{FID}}_{\infty}$ and $\overline{\text{KID}}_{\infty}$ scores, outperforming both MLE-trained KAEM and thermodynamic training. On SVHN, MLE-trained KAEM achieved metrics relatively close to VAE, while on CelebA the gap was more substantial. Visual inspection of samples in Appendix G.2 corroborates that VAE produced the most realistic samples on both datasets. However, the metrics otherwise did not reflect perceptual quality: on CelebA, thermodynamic training produced samples that appeared subjectively closer to VAE quality than MLE-trained KAEM, despite achieving worse scores.
This suggests that improved posterior sampling may be key to closing the gap with VAEs. Future work could investigate priors parameterized by functions other than wavelets, as well as more sophisticated posterior sampling strategies.
@@
Ultimately, we cannot justify the thermodynamic strategy due to its substantial computational overhead. With greater VRAM availability, it would be possible to parallelize a larger number of temperatures than the $N_{t}=5$ used in this study, which could potentially improve sample quality, as suggested by the results of Calderhead and Girolami (2009). Alternatively, different posterior sampling approaches, such as other annealing schedules, more ULA steps, or the use of autoMALA (Biron-Lattes et al., 2024) to provide bias correction and temperature-adaptive step-size tuning, may further improve performance. Alternatively, amortized inference approaches similar to VAEs could also be explored, as the information bottleneck criterion has been previously derived for latent EBMs by Kong et al. (2021).
@@ 6 Conclusions
KAEM provides a framework for structured, interpretable generative modeling. Our results show that importance sampling enables efficient training on simple or low-dimensional datasets, while the Kolmogorov-Arnold Representation Theorem offers a basis for architectural design, allowing KAEM’s structure to be determined entirely by its latent space dimensionality. The inverse transform method provides clear advantages over Langevin Monte Carlo for sampling energy-based priors, offering both speed and exactness. VAEs achieved the best image metrics on both SVHN and CelebA, with subjective inspection of samples in Appendix G.2 corroborating this assessment. However, thermodynamic training produced samples that appeared subjectively closer to VAE quality than MLE-trained KAEM, suggesting that improved posterior sampling may be a promising direction for closing the quality gap.
@@ 6.1 Future work
Moreover, while ITS enables fast and scalable inference for mixture priors, KAEM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of Langevin dynamics. GPUs are poorly suited to Langevin methods because of irregular memory access patterns and bandwidth pressure from differentiating with respect to inputs, which requires storing intermediate outputs across generator layers.
@@
Through other experiments, we found that the choice of prior architecture and its encoded inductive biases significantly affect sample quality. Future work could investigate this across different KAN bases and datasets, as well as explore how domain knowledge can guide the selection of univariate functions.
@@ F.1 Hyperparameters tables
generator_noise $\sigma_{\epsilon}$ 0.1
generator_variance $\sigma_{\epsilon}$ 0.3
@@
update_grid (-) false
Mixture
use_mixture_prior (-) true
train_proportions (-) true
mixture_regularization $\lambda$ 0.0001
@@
Number of epochs $N_{\text{epochs}}$ 100
eps $\epsilon$ 0.0001
@@ F.2.1 KAEM Architectures
Input $\bm{z}\to$ Dense(Q,2048)(Q,2048) $\to$ reshape $4\times 4\times 128$
1 $\times4$ convT($128$), ReLU $8\times 8\times 128$ 2
2 $\times4$ convT($64$), ReLU $16\times 16\times 64$ 2
@@ 3 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
Config: hidden_feature_dims=128,64, kernel=4,$4,$ strides=2/2/2, paddings=1/1/1, activation=ReLU, batchnorm=False, projection=True
@@
Input $\bm{z}\to$ Dense(Q,4096)(Q,4096) $\to$ reshape $4\times 4\times 256$
1 $\times4$ convT($256$), ReLU $8\times 8\times 256$ 2
2 $\times4$ convT($128$), ReLU $16\times 16\times 128$ 2
3 $\times4$ convT($64$), ReLU $32\times 32\times 64$ 2
@@ 4 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
Config: hidden_feature_dims=256,128,64, kernel=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, activation=ReLU, batchnorm=False, projection=True
@@ F.2.2 Baseline VAE architectures
1 $\times4$ conv($32$), ReLU $16\times 16\times 32$ 2
2 $\times4$ conv($64$), ReLU $8\times 8\times 64$ 2
3 $\times4$ conv($128$), ReLU $4\times 4\times 128$ 2
4 Flatten + Dense $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
Config: channels=32,64,128, kernels=4,$4,$ strides=2/2/2, paddings=1/1/1, latent_dim=81, batchnorm=False
@@ 4 Flatten + Dense $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
1 Dense(81,2048)(81,2048) + Reshape $4\times 4\times 128$
2 $\times4$ convT($128$), ReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), ReLU $16\times 16\times 64$ 2
@@ 4 $\times4$ convT(33), Sigmoid $32\times 32\times 3$ 2
Config: channels=128,64,32, kernels=4,$4,$ strides=2/2/2, paddings=1/1/1, latent_dim=81, batchnorm=False
@@
1 $\times4$ conv($32$), ReLU $32\times 32\times 32$ 2
2 $\times4$ conv($64$), ReLU $16\times 16\times 64$ 2
3 $\times4$ conv($128$), ReLU $8\times 8\times 128$ 2
4 $\times4$ conv($256$), ReLU $4\times 4\times 256$ 2
5 Flatten + Dense $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
Config: channels=32,64,128,256, kernels=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, latent_dim=81, batchnorm=False
@@ 5 Flatten + Dense $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
1 Dense(81,4096)(81,4096) + Reshape $4\times 4\times 256$
2 $\times4$ convT($256$), ReLU $8\times 8\times 256$ 2
3 $\times4$ convT($128$), ReLU $16\times 16\times 128$ 2
4 $\times4$ convT($64$), ReLU $32\times 32\times 64$ 2
@@ 5 $\times4$ convT(33), Sigmoid $64\times 64\times 3$ 2
Config: channels=256,128,64,32, kernels=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, latent_dim=81, batchnorm=False